a) A driver was speeding with 40 m/s in his 1800 kg car when he notices a police car with a radar and slows down to a legal speed of 19 m/s. What is the change of the kinetic energy of the car? What is the amount of work done on the car during the deceleration?


b) A repairman of mass 80 kg climbed a 10 m vertical latter to reach the roof of a building in 12 seconds. What is the power developed by the man against gravity?


1
Expert's answer
2022-04-11T12:03:46-0400

a)

"∆E=\frac m2(v_2^2-v_1^2)=-1.12~MJ,"

"A=-\u2206E=1.12~MJ."

b)

"P=\frac{mgh}t=654~W."